OpenStudy (anonymous):

evaluate: log base 32 to the 2

OpenStudy (amistre64):

logs are exponents

OpenStudy (amistre64):

\[log_{32}(2)=n\] \[32^{log_{32}(2)}=32^n\] \[2=32^n\]

OpenStudy (amistre64):

we can try a change of base, if you cant use a calculator on it\[log_{32}(2)=\frac{log_2(2)}{log_2(32)}=\frac{1}{?}\]

OpenStudy (amistre64):

2.2=4.2=8.2=16.2=32 ^ ^ ^ ^ ^ ; got 5 tics so im going to go with; 1/5

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

1/5?

OpenStudy (amistre64):

log2(32) = 5, so yes. the answer, if im reading the post correctly, should be 1/5

OpenStudy (anonymous):

ok thanks! because i wasnt there the day we went over this so i had no idea what to do or even were to start.

OpenStudy (amistre64):

if you want to get better at these, there are just a few basic log rules that you can practice.

OpenStudy (anonymous):

ok wha are they? love to hear em